OpenStudy (anonymous):

Which equivalence factor set should you use to convert 2.68 x 1011 atoms of silver (Ag) to grams of silver? (2.68 x 1011 atoms Ag/1 mol Ag)(1 mol Ag/107.88 g Ag) (1 mol Ag/2.68 x 1011 atoms Ag)(107.88 g Ag/1 mol Ag) (1 mol Ag/6.02 x 1023 atoms Ag)(107.88 g Ag/1 mol Ag) (2.68 x 1011 atoms Ag/6.02 x 1023 atoms Ag)(107.88 g Ag)

OpenStudy (whpalmer4):

set it up with the units and it will be easy. you have atoms, you want grams. You should know how to convert between atoms and moles, and between grams and moles. \[\text{atoms} * \frac{\text{mole}}{\text{atoms}} * \frac{\text{grams}}{\text{mole}} \rightarrow \cancel{\text{atoms}} *\frac{\cancel{\text{mole }}} {\cancel{\text{atoms}}} * \frac{{\text{grams }}} {\cancel{\text{mole}}}\rightarrow\text{grams}\]